OpenStudy (anonymous):

Find the product. Express the result in rectangular form. 4(cos(pi/3 + isin(pi/3) * 7(cos(2pi/3) + isin(2pi/3)

OpenStudy (anonymous):

multiply the numbers out front, add the angles that is all

OpenStudy (anonymous):

\[4\times 7=?\] \[\frac{\pi}{3}+\frac{2\pi}{3}=?\]

OpenStudy (anonymous):

28, pi

OpenStudy (jdoe0001):

as mentioned => \(\bf r_1[cos(\alpha)+isin(\alpha)]\cdot r_2[cos(\beta)+isin(\beta)] \\ \quad \\ \implies r_1\cdot r_2[cos(\alpha + \beta)+isin(\alpha + \beta)]\)

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

ok thnx cx

OpenStudy (anonymous):

\[28(\cos(\pi)+i\sin(\pi))\]

OpenStudy (anonymous):

it is ridiculously easy, which is pretty much the whole point of writing a complex number in trig form